OpenStudy (anonymous):

The area of a circle whose circumference is 8π is___π square units

jimthompson5910 (jim_thompson5910):

C = 2*pi*r 8pi = 2*pi*r Solve this for r

jimthompson5910 (jim_thompson5910):

What do you get?

OpenStudy (anonymous):

so would it be 4pi

OpenStudy (anonymous):

is that right

jimthompson5910 (jim_thompson5910):

no it should be just 4

jimthompson5910 (jim_thompson5910):

with no "pi" at the end of it